((1/2x) + 1/2) = (4/(x + 3))
Answer by stanbon(75887)

((1/2x) + 1/2) = (4/(x + 3))
---
(2+2x)/(4x) = 4/(x+3)
16x = (2+2x)(x+3)
16x = 2x+6 + 2x^2+6x
---
16x = 2x^2+8x+6
x^2+4x+3 = 8x
---
x^2-4x+3 = 0
(x-3)(x-1) = 0
x = 3 or x = 1
===================
Cheers,
Stan H.